OpenStudy (anonymous):

For each table, determine whether the relationship is a linear function. Then represent the relationship using an equation. 1. X Y 0 52.07 1 53.36 2 54.65 3 55.94 2. X Y 0 11.2 17 10.2 34 9.2 51 8.2 Please Help Medals for correct Answer!

OpenStudy (danjs):

linear will always change the same amount per change in x,

OpenStudy (danjs):

slope of th eline

OpenStudy (anonymous):

what they said ^^^

OpenStudy (danjs):

Check if the slope stays constant... for any 2 points in the table, the slope will be the same...

OpenStudy (anonymous):

thanks

OpenStudy (danjs):

if it is a linear set of points, then use any of the two points , to get a slope, and then construct the point-slope form for the line \[\large (y-y_{1})=m*(x-x _{1})\] for any point (x1,y1) in the table

OpenStudy (danjs):

Or since they give you the y-interept in the table (the b, in y=mx+b ) form of a line. just use that and the slope